OpenStudy (anonymous):

Calculus early transcendental functions fourth edition. f(x)= ln ((x^2)+1), c=0 find the Taylor Series. Can anyone please help me on this?

OpenStudy (anonymous):

sure you want to do it step by step?

OpenStudy (anonymous):

yes that would help me understand it so much!

OpenStudy (anonymous):

ok there is probably a snappy way to do it, but we can work it out slowly. you are expanding about 0 so first term is \[f(0)=\ln(1)=0\]

OpenStudy (anonymous):

isn't it 0?

OpenStudy (anonymous):

i am lost already what happened to the (x^2 +1)

OpenStudy (anonymous):

ok back up

OpenStudy (anonymous):

first term is \[f(0)\] \[f(x)=\ln(x^2+1)\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so \[f(0)=\ln(0^2+1)=\ln(1)=0\] is that ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now we need \[f'(0)\] because the next term will be \[f'(0)x\] and to find \[f'(0)\] we first find \[f'(x)\] then replace x by 0

OpenStudy (anonymous):

so our next job is to find the derivative of \[\ln(x^2+1)\] and this requires the chain rule. we get \[f'(x)=\frac{2x}{x^2+1}\]

OpenStudy (anonymous):

\[f'(x) = \frac{1}{x^2 + 1}* 2x = \frac{2x}{x^2+1}\]

OpenStudy (anonymous):

and therefore \[f'(0)=\frac{2\times 0}{0^2+1}=\frac{0}{1}=0\]

OpenStudy (anonymous):

so first two terms of the taylor series are 0

OpenStudy (anonymous):

\[f'(0) = 0\ => f'(0)*x = 0\]

OpenStudy (anonymous):

the next term will be \[\frac{f''(0)}{2}x^2\]

OpenStudy (anonymous):

and before we go any further, i give you a hint. \[\ln(x^2+1)\] is an even function, because if you replace x by -x you get the same thing back. so all terms will have only even powers. the odd powers will all have coefficient zero

OpenStudy (anonymous):

ok i understand the first derivative, n ot to sure about getting the second derivative. I get the even function so no odd exponents

OpenStudy (anonymous):

doctor agdgdgdgwngo has found the second derivative. but i think there is a mistake there

OpenStudy (anonymous):

lets check hold on

OpenStudy (anonymous):

There is: it should be \frac{-2(x^2-1)}{(x^2+1)^2}

OpenStudy (anonymous):

how did you get the second derivative?

OpenStudy (anonymous):

\[f''(x) = \frac{-2(x^2-1)}{(x^2+1)^2}\]

OpenStudy (anonymous):

you take the derivative of the first derivative

OpenStudy (anonymous):

\[f''(0) = 2\] \[\frac{f''(0)}{2}x^2 = x^2\]

OpenStudy (anonymous):

now \[f''(0)=2\] so \[\frac{f''(0)}{2}x^2=x^2\] that is the first term of the series

OpenStudy (anonymous):

next term will be 0 because you have no odd powers here

OpenStudy (anonymous):

yes I understand that I was just wondering what rule you used to get the second derivative

OpenStudy (anonymous):

quotient probably right?

OpenStudy (anonymous):

there is a much easier way to do this, because these derivatives are really going to suck in a minute, but the answer is very simple

OpenStudy (anonymous):

i dont get that the equation, (second derivative(0)/2 )=x^2

OpenStudy (anonymous):

third derivative = \[f'''(x)=\frac{4x(x^2-3)}{(x^2+1)^3}\] fourth derivative=\[f''''(x) = \frac{-12(x^4-6x^2+1)}{(x^2+1)^4}\]

OpenStudy (anonymous):

yes it is \[\frac{f''(0)}{2}x^2\] and a general terms looks like \[\frac{f^{(n)}(0)}{n!}x^n\]

OpenStudy (anonymous):

lets please do it the easy way ok?

OpenStudy (anonymous):

fifth derivative\[f'''''(x) = \frac{48x(x^4-10x^2+5)}{(x^2+1)^5}\]

OpenStudy (anonymous):

tell me if I am getting this, you keep on taking the derivatives and plugging in 0 for x then because you were given c=0?

OpenStudy (anonymous):

because these derivatives really get annoying fast

OpenStudy (anonymous):

ok satellite I get that formula

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can we do it the easy way now please?

OpenStudy (anonymous):

yes go ahead

OpenStudy (anonymous):

instead of finding the expansion for \[f(x)=\ln(x^2+1)\] instead find it for \[\ln(x+1)\] which is both well known and much much easier to find. ready?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the derivatives are really simple, you do not need the quotient rule

OpenStudy (anonymous):

\[f'(x)=\frac{1}{x+1}=(x+1)^{-1}\] \[f''(x)=-1(x+1)^{-2}\] \[f'''(x)=2(x+1)^{-3}\] \[f^{(4)}(x)=-6(x+1)^{-4}\] etc

OpenStudy (anonymous):

I smell factorial

OpenStudy (anonymous):

ok i get how you got those derivatives

OpenStudy (anonymous):

and in general from the pattern you see \[f^{(n)}(x)=(-1)^{n-1}(n-1)!(x+1)^{-n}\]

OpenStudy (anonymous):

now we only want these at x = 0 so really we are only interested in the coefficients right?

OpenStudy (anonymous):

for example \[f'(0)=1\] \[f''(0)=-1\] \[f'''(0)=2\] \[f^{(4)}(0)=-6\] \[f^{(5)}(0)=4!\] and so on

OpenStudy (anonymous):

whew. and now since we are dividing each of these by n! we have each term looks like \[(-1)^{n-1}\frac{x^n}{n}\]

OpenStudy (anonymous):

this because \[\frac{(n-1)!}{n!}=\frac{1}{n}\]

OpenStudy (anonymous):

let me know if i have completely lost you and we can back up

OpenStudy (anonymous):

i got -1 when I plugged in n=1 and x=0 into f prime but understand what u are doing, i just dont know where you got that formula, and not getting we are dividing by n! now part

OpenStudy (anonymous):

ok lets to back to the formula and worry about the n! later

OpenStudy (anonymous):

the derivatives look like \[(x+1)^{-1}\] \[-(x+1)^{-2}\] \[2(x+1)^{-3}\] \[-6(x+1)^{-4}\] \[24(x+1)^{-5}\] what would the next one be?

OpenStudy (anonymous):

120(x+1)^-6

OpenStudy (anonymous):

sorry -120(x+1)^-6

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

and since we want these at x = 0 we are really only interested in the numbers out front right? they are 1 -1 2 -6 24 -120 etc

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and where did the -120 come from? it came from -5*24 and the 24 came from 4*6 and the 6 came from 3*2 so we found 120 via \[120=6\times 5\times 4\times 3\times 2=6!\]

OpenStudy (anonymous):

so each time you take a derivative, you are multiplying a successive number by what you had previously

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

that is why the nth derivative has coefficient (n-1)!

OpenStudy (anonymous):

ok formula i will have to know thank you for explaining that to me

OpenStudy (anonymous):

well not exactly, we have to adjust for the fact that it is alternating. so the coefficient of the nth derivative is actually \[(-1)^{n-1}(n-1)!\]

OpenStudy (anonymous):

ok with the -1 it alternates from positive t negative right

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now we can finish more easily. each term looks like \[\frac{f^{(n)}(0)}{n!}x^n\]

OpenStudy (anonymous):

and we know \[f^{(n)}(0)=(-1)^{n-1}(n-1)!\]

OpenStudy (anonymous):

and we also know that \[\frac{(n-1)!}{n!}=\frac{1}{n}\]

OpenStudy (anonymous):

so each term of the expansion of \[\ln(x+1)\] is \[(-1)^{n-1}\frac{x^n}{n}\]

OpenStudy (anonymous):

in other words \[\ln(x+1)=x-\frac{x^2}{2}+\frac{x^3}{3}-\frac{x^4}{4} +...\] it is a simple one

OpenStudy (anonymous):

also a well known one. you see the pattern immediately. and finally (finally) we are ready to solve your problem because you did not have \[\ln(x+1)\] you had \[\ln(x^2+1)\]

OpenStudy (anonymous):

ok how do we get from there to (x^2+1)

OpenStudy (anonymous):

ok well all the hard work was done, even though the answer is fairly simple. now since you know the expansion for \[\ln(x+1)\] if you want the expansion for \[\ln(x^2+1)\] you ... replace x by x^2

OpenStudy (anonymous):

and get \[x^2-\frac{x^4}{2}+\frac{x^6}{3}-\frac{x^8}{4}+\frac{x^{10}}{5}-...\]

OpenStudy (anonymous):

if you want fancy notation each term looks like \[(-1)^{k-1}\frac{x^{2k}}{k}\]

OpenStudy (anonymous):

ok and if I had lnx and knew that answer and was asked lnx^2 i would do the same thing?

OpenStudy (anonymous):

so you can write \[\ln(x^2+1)=\sum_{k=1}^{\infty}\frac{(-1)^{k-1}x^{2k}}{k}\]

OpenStudy (anonymous):

well you cannot expand \[\ln(x)\] about 0 because \[\ln(0)\] is undefined that is why you generally see it as the expansion of \[\ln(x+1)\] about 0

OpenStudy (anonymous):

you can expand \[\ln(x)\] about 1 if you like

OpenStudy (anonymous):

you get the same thing we got with \[\ln(x+1)\] just with \[(x-1)\] in place of x

OpenStudy (anonymous):

yes thats the formula i have, thank you so much for explaining this to me,taking online calc 2 course and I spend so much time looking elsewhere trying to anwer hw cause i can't get the answers out of this book thank you so much

OpenStudy (anonymous):

\[\ln(x)=(x-1)-\frac{(x-1)^2}{2}+\frac{(x-1)^3}{3}-...\]

OpenStudy (anonymous):

yw and good luck. btw if it is homework on line cheat when you can

OpenStudy (anonymous):

i try, sometimes its the only way that I can understand how to do the problems, I am self teaching myself calc 2 whith a bad book instead of reading and studying I waste an unbelievable amount of time trying to find answers and explainations to this course